Session: Feb 1998 (session 1)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (GMaccaferri at Sat Feb 21 19:22:47 MET 1998) Success

Onsala (A.Polatidis at Sat Feb 21 20:38:38 MET 1998) Success

Torun (K.Borkowski at Mon Feb 23 10:04:29 MET 1998) Success

Westerbork (Tony Foley at Mon Feb 23 10:52:43 MET 1998) Success
Looks fine. Some EMI noticed at start. Parity errors returned are wrong as the standard procedures used the wrong tape speed.

Jodrell (Alastair Gunn at Mon Feb 23 11:17:42 MET 1998) Minor failures
Extreme RFI in VCs 9 and 10. Difficulty maintaining appropriate attenuation levels. No other known problems.

Noto (C.Trigilio at Wed Feb 25 14:27:58 MET 1998) Minor failures
BBC08 unlocked for all the experiment.

Urumqi (Liu Xiang at Fri Feb 27 04:58:22 MET 1998) Success

Effelsberg (dgraham at Thu Mar 5 16:00:05 MET 1998) Success

JIVE (Lorant Sjouwerman at Tue Apr 7 16:59:14 METDST 1998) Success
Network monitoring report of L-band session